sum of two unit vector is also a  unit vector then maganitude of their difference is

Grade:11

2 Answers

Priyanshu Gujjar
2 years ago

a+b=1

a2+∣b2+2∣a∣∣b∣cosθ=1
2(1+cosθ)=1
cosθ=21θ=120
a+b=a∣+∣b2−2∣a∣∣b∣cosθ
=1+1−2cos120
=2+2×21=3.
Heer Shah
2 years ago
let a,b be two unit vector at angle q
And c be there resultant unit vector.
Thus 1=a²+b²+2abcos(q)
1=1+1+2.1.1.cos(q)
cosq= -1/2
q=2pi/3
 
|a-b|=root(a²+b²+2ab cos(pi-2pi/3)
(Pi-2pi/3) because direction of b is reverse and we are adding (-b) to A vector.
|a-b|=root(1+1+2×1/2)=root(3)